OpenStudy (anonymous):

What is 8(4k-4)=5k-32

OpenStudy (akashdeepdeb):

Use the distributive property of arithmetic, open the brackets, thus. 8(4k-4) = 5k-32 32k - 32 = 5k-32 Thus 32k = 5k k = 0 !!! :D